7.) State the Fundamental Theorem of Arithmetic and use it to prove that
3 p
625 is irrational.7.) State the Fundamental Theorem of Arithmetic and use it to prove that 625 is irrational.
Answer #1

\underline{FUNDAMENTAL\; THEOREM \;OF \;ARITHMETIC :-}\\ According \;to \;the\; fundamental \;theorem\; of\; arithmetic\; every \;composite\; number \;\\can\; be\; written\; or \; factorized\; as \;the\; product\; of\; primes\; in\; an\; unique\; way\; apart\; \\from\; the \;order\; in\; which\; the\; prime\; factors\; occur.\\ It\; is\; also\; called\; as\; UNIQUE \;FACTORIZATION\; THEOREM.\\ \\ Composite\; number =\; product \;of\; prime\; numbers\\ \\\; Or\\ Any\; integer\; greater\; than\; one\; can\; either\; be\; a \;prime \;number\; or\; can\; be \;written\; as\\ \;a \;product\; of\; prime\; factors\;uniquely.

To\; prove:-\; \sqrt[3]{625}\; is\; irrational.\\Proof:-\\ we\; prove\; it\; by\; the\; method\; of\; contradiction.\\Let\; \sqrt[3]{625}\; is\; a\; rational\; number.\\So,\; \sqrt[3]{625}=\frac{a}{b},\;\;\;where\;a\;and\;b\;are\; co-prime\; integers\;(i.e.,\; HCF\; of\; a\; and\; b\; is\;1).\\ Cubing\; both\; the\; sides\; we\; get\\ 625=\frac{a^3}{b^3}\\ \implies 625b^3=a^3\;\;\;\;\;\;-----(i)\\ Now\; by\; Fundamental\; Theorem\; of\; Arithmetic,\; this\; implies\; a^3\; is\; divisible\; by\; 625\; \\and\; subsequently\; a\; is\; divisible\; by\;625\\i.e.,\; 625|a^3 \implies 625|a\;\;\;\;\;\;-----(ii)\\ Then\;\exists \;k\in\mathbb{Z},\; such\; that\; a=625k.\\From\;(i)\\ 625b^3=a^3\implies 625b^3=(625k)^3\implies 625b^3=625^3a^3\implies b^3=625^2a^3\\ Now\;again\; by\; Fundamental\; Theorem\; of\; Arithmetic,\; this\; implies\; b^3\; is\; divisible\; by\; 625\; \\and\; subsequently\; b\; is\; divisible\; by\;625\\i.e.,\; 625|b^3 \implies 625|b\;\;\;\;\;\;-----(iii)\\So,\; from\; (ii)\;and\;(iii)\; we\; obtain\; that\; 625\; is\; a\; common\; factor\; of\; 'a'\;and\;'b'\;.\\But\; this\; contradicts\; the\; fact\; that\; a\; and\; b\; are\; co-prime\;intgers.\\So\; our\; assumption\; of\; \sqrt[3]{625}\; is\; a\; rational\; is\; wrong.\\Hence,\;\;\sqrt[3]{625}\; is\; an\; irrational\; number.\\\;\\\;\\\;\\Note:- \\As\; \sqrt[3]{625}=5\sqrt[3]{5}\; so,\; you\; may\; also\; prove\; \sqrt[3]{625}\;as\; irrational\; by\; proving\; \sqrt[3]{5}\; irrational\;\\ by\; the\; same\; method\; of \; contradiction\; as\;I\; proved \;above.